Calculate the wavelength of light absorb when an electron jump from n=5 to n=6 state in hydrogen atom?

1
Expert's answer
2022-01-10T13:56:50-0500

"E_n=-2.18×10^{-18}J(\frac{1}{n^2})"


"=-2.18×10^{-18}J(\frac{1}{n^2_f}-\frac{1}{n^2_i})"


"=-2.18×10^{-18}J(\frac{1}{6^2}-\frac{1}{5^2})"


="-2.18×10^{-18}J(-\frac{11}{900})"


"=2.66\u00d710^{-20}J"


"/\Delta E / =E_{photon}=hv =\frac{hc}{\lambda}"

Where h is Planck's constant,

c is the speed of light,

"\lambda" is the wavelength of the incoming photon

Thus wavelength is

"\lambda=\frac{hc}{E_{photon}}"


"=\frac{(6.626×10^{-34}J.s)(2.998×10^8m/s)}{2.66×10^{-20}J}"


"=7.468\u00d710^{-6}m"


"=74.68nm"